A stone is thrown vertically upward with a speed of 18.0m/s. How fast is it moving when it reaches a height of 11.0m? How long is required to reach this height?Why are there two answers?
Answer #1
h =ut -0.5 gt^2
or, 11 = 18t -4.9 t^2
t =0.774 sec

v = 18-9.8*.774 =10.418 m/s
answered by: CSI